If the nominal rate of interest is

13.94 %13.94%

and the real rate of interest is

7.3 %7.3%​,

what is the expected rate of​ inflation?

The expected rate of inflation is

nothing​%
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.073)=(1+0.1394)/(1+inflation rate)

1+inflation rate=(1.1394/1.073)

inflation rate=(1.1394/1.073)-1

=6.19%(Approx).